1.     Assume a hypothetical consumer consumes good X and good Y. The price of good X is 1 and price of good Y is 3 and the consumer budget is birr 10 for the two goods. Where: QX is quantity of good X, QY is quantity of good Y and TUX and TUY is total utility from consuming good X and good Y respectively.

What are the major properties of indifference curves?

 

                   Express the budget line of the consumer both algebraically and diagrammatically


1
Expert's answer
2021-04-19T07:42:15-0400

Solution:

The major properties of indifference curves are the following:

·        The indifference curves slope downwards.

·        The indifference curves are convex towards the origin.

·        Higher indifference curves represent higher utility and are more preferred.

·        Indifference curves cannot intersect.

 

Budget line algebraically:

I = Px*QX + Py*QY

I = PxQX + PyQY

10 = QX + 3QY

 

Budget diagrammatically:

Good X qty = 10/1 = 10

Good Y qty = 10/3 = 3.33

See below graph: